OpenStudy (anonymous):

simplify the following surd 6√64

OpenStudy (anonymous):

6 x 8 = 48

OpenStudy (anonymous):

what happens to the 64 though?

OpenStudy (anonymous):

64 is 8x8 so 8^2

OpenStudy (anonymous):

It becomes 8. What it's asking is basically what number times itself is 64? and that number happens to be 8. 8x8 = 64

OpenStudy (anonymous):

oh okay, cheers.

OpenStudy (anonymous):

how come 7√50 is different?

OpenStudy (anonymous):

it equals 35√2

OpenStudy (lalaly):

\[7\sqrt{50}\]\[\sqrt50= \sqrt{25} \times \sqrt{2}\]and\[\sqrt25 = 5\]so \[7\sqrt50 = 7 \times \sqrt25 \times \sqrt2\]\[=7 \times 5 \times \sqrt2\]\[=35\sqrt2\]